# AI Risk Atlas: Taxonomy and Tooling for Navigating AI Risks and Resources

### v1

@@ 2 Conclusion and call to action
Prompt injection attack A prompt injection attack forces a generative model that takes a prompt as input to produce unexpected output by manipulating the structure, instructions, or information contained in its prompt.
@@
Type: output
Descriptor: specific
Implementation details:
ID: atlas-incomplete-advice
Tag: incomplete-advice
URI: IBM AI Risk Atlas - Incomplete advice
@@
Extraction attack An attribute inference attack is used to detect whether certain sensitive features can be inferred about individuals who participated in training a model. These attacks occur when an adversary has some prior knowledge about the training data and uses that knowledge to infer the sensitive data.
@@
Type: non-technical
Descriptor: amplified
Implementation details:
ID: atlas-job-loss
Tag: job-loss
URI: IBM AI Risk Atlas - Impact on Jobs
Jailbreaking A jailbreaking attack attempts to break through the guardrails that are established in the model to perform restricted actions.
Concern: Jailbreaking attacks can be used to alter model behavior and benefit the attacker. If not properly controlled, business entities can face fines, reputational harm, and other legal consequences.
@@
Concern: ”There are several ways of collecting data for building a foundation models: web scraping, web crawling, crowdsourcing, and curating public datasets. Data acquisition restrictions can also impact the availability of the data that is required for training an AI model and can lead to poorly represented data.”
@@
Type: training-data
Descriptor: amplified
@@
ID: atlas-data-bias
Tag: data-bias
URI: IBM AI Risk Atlas - Data bias
Type: training-data
Descriptor: amplified
Implementation details:
ID: atlas-data-provenance
Tag: data-provenance
URI: IBM AI Risk Atlas - Uncertain data provenance
Type: non-technical
Descriptor: amplified
Implementation details:
ID: atlas-unrepresentative-risk-testing
Tag: unrepresentative-risk-testing
URI: IBM AI Risk Atlas - Unrepresentative risk testing
Type: training-data
Descriptor: amplified
Implementation details:
ID: atlas-data-usage-rights
Tag: data-usage-rights
URI: IBM AI Risk Atlas - Data usage rights restrictions
Type: output
Descriptor: specific
Implementation details:
ID: atlas-harmful-code-generation
Tag: harmful-code-generation
URI: IBM AI Risk Atlas - Harmful code generation
Type: training-data
Descriptor: amplified
Implementation details:
ID: atlas-data-contamination
Tag: data-contamination
URI: IBM AI Risk Atlas - Data contamination
Type: non-technical
Descriptor: specific
Implementation details:
ID: atlas-incomplete-usage-definition
Tag: incomplete-usage-definition
URI: IBM AI Risk Atlas - Incomplete usage definition
Type: output
Descriptor: specific
Implementation details:
ID: atlas-copyright-infringement
Tag: copyright-infringement
URI: IBM AI Risk Atlas - Copyright infringement
Type: non-technical
Descriptor: amplified
Implementation details:
ID: atlas-lack-of-data-transparency
Tag: lack-of-data-transparency
URI: IBM AI Risk Atlas - Lack of data transparency
Type: non-technical
Descriptor: traditional
Implementation details:
ID: atlas-impact-on-affected-communities
Tag: impact-on-affected-communities
URI: IBM AI Risk Atlas - Impact on affected communities
Type: training-data
Descriptor: amplified
Implementation details:
ID: atlas-improper-retraining
Tag: improper-retraining
URI: IBM AI Risk Atlas - Improper retraining
Type: output
Descriptor: specific
Implementation details:
ID: atlas-spreading-toxicity
Tag: spreading-toxicity
URI: IBM AI Risk Atlas - Spreading toxicity
Type: output
Descriptor: amplified
Implementation details:
ID: atlas-inaccessible-training-data
Tag: inaccessible-training-data
URI: IBM AI Risk Atlas - Inaccessible training data
Type: non-technical
Descriptor: specific
Implementation details:
ID: atlas-bypassing-learning
Tag: bypassing-learning
URI: IBM AI Risk Atlas - Impact on education: bypassing learning
Type: output
Descriptor: amplified
Implementation details:
ID: atlas-untraceable-attribution
Tag: untraceable-attribution
URI: IBM AI Risk Atlas - Untraceable attribution
Evasion attack Evasion attacks attempt to make a model output incorrect results by slightly perturbing the input data that is sent to the trained model.
@@
Descriptor: amplified
Implementation details:
ID: atlas-evasion-attack
Tag: evasion-attack
URI: IBM AI Risk Atlas - Evasion attack
Type: non-technical
Descriptor: amplified
Implementation details:
ID: atlas-impact-on-the-environment
Tag: impact-on-the-environment
URI: IBM AI Risk Atlas - Impact on the environment
Type: output
Descriptor: amplified
Implementation details:
ID: atlas-over-or-under-reliance
Tag: over-or-under-reliance
URI: IBM AI Risk Atlas - Over- or under-reliance
Type: non-technical
Descriptor: amplified
Implementation details:
ID: atlas-incorrect-risk-testing
Tag: incorrect-risk-testing
URI: IBM AI Risk Atlas - Incorrect risk testing
Membership inference attack A membership inference attack repeatedly queries a model to determine whether a given input was part of the model’s training. More specifically, given a trained model and a data sample, an attacker samples the input space, observing outputs to deduce whether that sample was part of the model’s training.
Concern: Identifying whether a data sample was used for training data can reveal what data was used to train a model. Possibly giving competitors insight into how a model was trained and the opportunity to replicate the model or tamper with it. Models that include publicly-available data are at higher risk of such attacks.
Type: inference
Descriptor: amplified
Implementation details:
ID: atlas-membership-inference-attack
Tag: membership-inference-attack
URI: IBM AI Risk Atlas - Membership inference attack
Type: inference
Descriptor: specific
Implementation details:
ID: atlas-confidential-data-in-prompt
Tag: confidential-data-in-prompt
URI: IBM AI Risk Atlas - Confidential data in prompt
Type: training-data
Descriptor: amplified
Implementation details:
ID: atlas-data-privacy-rights
Tag: data-privacy-rights
URI: IBM AI Risk Atlas - Data privacy rights alignment
Type: inference
Descriptor: specific
Implementation details:
ID: atlas-ip-information-in-prompt
Tag: ip-information-in-prompt
URI: IBM AI Risk Atlas - IP information in prompt
Concern: A successful attack copies the system prompt used in the model. Depending on the content of that prompt, the attacker might gain access to valuable information, such as sensitive personal information or intellectual property, and might be able to replicate some of the functionality of the model.
Type: inference
Descriptor: specific
Implementation details:
ID: atlas-prompt-leaking
Tag: prompt-leaking
URI: IBM AI Risk Atlas - Prompt leaking
Type: output
Descriptor: specific
Implementation details:
ID: atlas-hallucination
Tag: hallucination
URI: IBM AI Risk Atlas - Hallucination
Type: non-technical
Descriptor: amplified
Implementation details:
ID: atlas-legal-accountability
Tag: legal-accountability
URI: IBM AI Risk Atlas - Legal accountability
Prompt priming Because generative models tend to produce output like the input provided, the model can be prompted to reveal specific kinds of information. For example, adding personal information in the prompt increases its likelihood of generating similar kinds of personal information in its output. If personal data was included as part of the model’s training, there is a possibility it could be revealed.
Type: inference
Descriptor: specific

### v2

@@ Appendix A AI Risk Atlas Definitions
Evasion attack Evasion attacks attempt to make a model output incorrect results by slightly perturbing the input data sent to the trained model.
Type: inference
Descriptor: amplified by generative AI
Implementation details:
ID: atlas-evasion-attack
Tag: evasion-attack
URI: IBM AI Risk Atlas - Evasion attack
Type: non-technical
Descriptor: amplified by generative AI
Implementation details:
ID: atlas-impact-on-the-environment
Tag: impact-on-the-environment
URI: IBM AI Risk Atlas - Impact on the environment
Type: non-technical
Descriptor: amplified by generative AI
Implementation details:
ID: atlas-incorrect-risk-testing
Tag: incorrect-risk-testing
URI: IBM AI Risk Atlas - Incorrect risk testing
Type: output
Descriptor: amplified by generative AI
Implementation details:
ID: atlas-over-or-under-reliance
Tag: over-or-under-reliance
URI: IBM AI Risk Atlas - Over- or under-reliance
Membership inference attack A membership inference attack repeatedly queries a model to determine if a given input was part of the model’s training. More specifically, given a trained model and a data sample, an attacker appropriately samples the input space, observing outputs to deduce whether that sample was part of the model’s training.
Concern: Identifying whether a data sample was used for training data can reveal what data was used to train a model, possibly giving competitors insight into how a model was trained and the opportunity to replicate the model or tamper with it. Models that include publicly-available data are at higher risk of such attacks.
Type: inference
Descriptor: amplified by generative AI
Implementation details:
ID: atlas-membership-inference-attack
Tag: membership-inference-attack
URI: IBM AI Risk Atlas - Membership inference attack
Type: inference
Descriptor: specific to generative AI
Implementation details:
ID: atlas-confidential-data-in-prompt
Tag: confidential-data-in-prompt
URI: IBM AI Risk Atlas - Confidential data in prompt
Concern: A successful prompt leaking attack copies the system prompt used in the model. Depending on the content of that prompt, the attacker might gain access to valuable information, such as sensitive personal information or intellectual property, and might be able to replicate some of the functionality of the model.
Type: inference
Descriptor: specific to generative AI
Implementation details:
ID: atlas-prompt-leaking
Tag: prompt-leaking
URI: IBM AI Risk Atlas - Prompt leaking
Type: training-data
Descriptor: amplified by generative AI
Implementation details:
ID: atlas-data-privacy-rights
Tag: data-privacy-rights
URI: IBM AI Risk Atlas - Data privacy rights alignment
Discriminatory actions AI agents can take actions where one group of humans is unfairly advantaged over another due to the decisions of the model. This may be caused by AI agents’ biased actions that impact the world, in the resources consulted, and in the resource selection process. For example, an AI agent can generate code that can be biased.
Concern: Discriminatory actions can cause harm to people. Discriminatory actions taken by an AI agent could perpetuate bias to systems outside the AI agent owner’s control, impact people, or lead to unintended consequences.
Type: agentic
Descriptor: amplified by agentic AI
Implementation details:
ID: atlas-discriminatory-actions
Tag: discriminatory-actions
URI: IBM AI Risk Atlas - Discriminatory actions
Type: inference
Descriptor: specific to generative AI
Implementation details:
ID: atlas-ip-information-in-prompt
Tag: ip-information-in-prompt
URI: IBM AI Risk Atlas - IP information in prompt
Type: non-technical
Descriptor: amplified by generative AI
Implementation details:
ID: atlas-legal-accountability
Tag: legal-accountability
URI: IBM AI Risk Atlas - Legal accountability
Type: output
Descriptor: specific to generative AI
Implementation details:
ID: atlas-hallucination
Tag: hallucination
URI: IBM AI Risk Atlas - Hallucination
Social hacking attack Manipulative prompts that use social engineering techniques, such as role-playing or hypothetical scenarios, to persuade the model into generating harmful content.
Concern: Social hacking attacks can be used to alter model behavior and benefit the attacker. The content it generates may cause harms for the user or others.
Type: inference
Descriptor: specific to generative AI
Implementation details:
ID: atlas-social-hacking-attack
Tag: social-hacking-attack
URI: IBM AI Risk Atlas - Social hacking attack
Harmful output A model might generate language that leads to physical harm. The language might include overtly violent, covertly dangerous, or otherwise indirectly unsafe statements.
Type: output
Descriptor: specific to generative AI
Implementation details:
ID: atlas-harmful-output
Tag: harmful-output
URI: IBM AI Risk Atlas - Harmful output
Indirect instructions attack Prompts, questions, or requests designed to elicit undesirable responses from the application. Unlike direct instructions attacks, the model is instructed to use instructions that are embedded in external data like a website.
Concern: Indirect instructions attacks can be used to alter model behavior and benefit the attacker. The content it generates may cause harms for the user or others.
Type: inference
Descriptor: specific to generative AI
Implementation details:
ID: atlas-indirect-instructions-attack
Tag: indirect-instructions-attack
URI: IBM AI Risk Atlas - Indirect instructions attack
Mitigation and maintenance The large number of components and dependencies that agent systems have complicates keeping them up to date and correcting problems.
Concern: AI agents may interact with other systems, tools, or other agents. Tracing the root cause of failure becomes more difficult and more costly as agent capabilities and complexities increase.
Type: agentic
Descriptor: amplified by agentic AI
Implementation details:
ID: atlas-mitigation-maintenance
Tag: mitigation-maintenance
URI: IBM AI Risk Atlas - Mitigation and maintenance
AI agent compliance Determining AI agents’ compliance is complex and there might not be enough information to assess whether the agentic AI system is compliant with applicable legal requirements.
Concern: AI agents may interact with other systems, tools, or other agents. AI agents can also find solutions to accomplish a task or a goal in a variety of ways and there could be uncertainty around the way an AI agent would choose each time to perform the task. Assessing compliance can become more difficult as agent capabilities increase.
Type: agentic
Descriptor: amplified by agentic AI
Implementation details:
ID: atlas-ai-agent-compliance
Tag: ai-agent-compliance
URI: IBM AI Risk Atlas - AI agent compliance
Function calling hallucination AI agents might make mistakes when generating function calls (calls to tools to execute actions). Those function calls might result in incorrect, unnecessary or harmful actions. Examples: Generating wrong functions or wrong parameters for the functions.
Concern: Hallucinations when generating function calls might result in wrong or redundant actions being performed. Depending on the actions taken, AI agents can cause harms to owners and users of the AI agents.
Type: agentic
Descriptor: specific to agentic AI
Implementation details:
ID: atlas-function-calling-hallucination
Tag: function-calling-hallucination
URI: IBM AI Risk Atlas - Function calling hallucination
Type: training-data
Descriptor: amplified by generative AI
Implementation details:
ID: atlas-confidential-information-in-data
Tag: confidential-information-in-data
URI: IBM AI Risk Atlas - Confidential information in data
Type: non-technical
Descriptor: traditional risk of AI
Implementation details:
ID: atlas-lack-of-model-transparency
Tag: lack-of-model-transparency
URI: IBM AI Risk Atlas - Lack of model transparency
Exploit trust mismatch Attackers might initiate injection attacks to bypass the trust boundary, which is a distinct point or conceptual line where the level of trust in a system, application or network changes. Background execution in multi-agent environments increases the risk of covert channels if input/output validation is weak.
Concern: This could lead to mismatched (expected vs. realized) trust boundaries and could result in unintended tool use, excessive agency, and privilege escalation.
Type: agentic
Descriptor: amplified by agentic AI
Implementation details:
ID: atlas-exploit-trust-mismatch
Tag: exploit-trust-mismatch
URI: IBM AI Risk Atlas - Exploit trust mismatch
Type: training-data
Descriptor: traditional risk of AI
Implementation details:
ID: atlas-unrepresentative-data
Tag: unrepresentative-data
URI: IBM AI Risk Atlas - Unrepresentative data
AI agents’ impact on human agency The autonomous nature of AI agents in performing tasks or taking actions could affect the individuals’ ability to engage in critical thinking, make choices and act independently.
Concern: AI agents might shift the decision, thinking, and control from humans to machines. This might negatively impact the society and human welfare as they limit the freedom and meaningful participations of humans in performing a task or making decisions.
Type: agentic
Descriptor: amplified by agentic AI
Implementation details:
ID: atlas-impact-human-agency
Tag: impact-human-agency
URI: IBM AI Risk Atlas - AI agents’ impact on human agency
Type: inference
Descriptor: specific to generative AI
Implementation details:
ID: atlas-personal-information-in-prompt
Tag: personal-information-in-prompt
URI: IBM AI Risk Atlas - Personal information in prompt
AI agents’ Impact on human agency The autonomous nature of AI agents in performing tasks or taking actions might affect the individuals’ ability to engage in critical thinking, make choices, and acting independently.
Concern: AI agents might shift the decision, thinking, and control from humans to machines. This might negatively impact society and human welfare as they limit the freedom and meaningful participations of humans in performing a task or making decisions.
Type: non-technical
Descriptor: amplified by generative AI
Implementation details:
ID: atlas-impact-on-human-agency
Tag: impact-on-human-agency
URI: IBM AI Risk Atlas - AI agents’ Impact on human agency
Sharing IP/PI/confidential information with user AI agents with unrestricted access to resources or databases or tools could potentially store and share PI/IP/confidential information with system users when performing their actions.
Concern: AI agents may share privileged information to users. The act of sharing the information may result in harm for the model owner, user, or others. The harm can vary based on the type and details of the information shared. Without adequate oversight, these privacy incidents might overwhelm company resources.
Type: agentic
Descriptor: amplified by agentic AI
Implementation details:
ID: atlas-sharing-info-user
Tag: sharing-info-user
URI: IBM AI Risk Atlas - Sharing IP/PI/confidential information with user
Type: non-technical
Descriptor: amplified by generative AI
Implementation details:
ID: atlas-lack-of-testing-diversity
Tag: lack-of-testing-diversity
URI: IBM AI Risk Atlas - Lack of testing diversity
Type: output
Descriptor: specific to generative AI
Implementation details:
ID: atlas-nonconsensual-use
Tag: nonconsensual-use
URI: IBM AI Risk Atlas - Nonconsensual use
Concern: Bias can harm persons affected by the decisions of the model.
Type: output
Descriptor: traditional risk of AI
Implementation details:
ID: atlas-decision-bias
Tag: decision-bias
URI: IBM AI Risk Atlas - Decision bias
Type: output
Descriptor: amplified by generative AI
Implementation details:
ID: atlas-exposing-personal-information
Tag: exposing-personal-information
URI: IBM AI Risk Atlas - Exposing personal information
AI agents’ impact on jobs Widespread adoption of AI agents to perform complex tasks might lead to widespread automation of roles and could lead to job displacement.
Concern: As trust in agentic systems increases, business may be more motivated to use agents instead of people. Job displacement might lead to a loss of income and thus might negatively impact society and human welfare. Re-skilling may be challenging given the pace of the technology evolution.
Type: agentic
Descriptor: amplified by agentic AI
Implementation details:
ID: atlas-impact-jobs
Tag: impact-jobs
URI: IBM AI Risk Atlas - AI agents’ impact on jobs
Type: training-data
Descriptor: amplified by generative AI
Implementation details:
ID: atlas-data-curation
Tag: data-curation
URI: IBM AI Risk Atlas - Improper data curation
Over- or under-reliance on AI agents Reliance, that is the willingness to accept an AI agent behavior, depends on how much a user trusts that agent and what they are using it for. Over-reliance occurs when a user puts too much trust in an AI agent, accepting an AI agent’s behavior even when it is likely undesired. Under-reliance is the opposite, where the user doesn’t trust the AI agent but should. Increasing autonomy (to take action, select and consult resources/tools) of AI agents and the possibility of opaqueness and open-endedness increase the variability and visibility of agent behavior leading to difficulty in calibrating trust and possibly contributing to both over- and under-reliance.
Concern: Over/under reliance can lead to poor decision making by humans because of their misplaced trust in the AI agent, with negative consequences that escalate with the significance of the decision.
Type: agentic
Descriptor: amplified by agentic AI
Implementation details:
ID: atlas-over-or-under-reliance-on-ai-agents
Tag: over-or-under-reliance-on-ai-agents
URI: IBM AI Risk Atlas - Over- or under-reliance on AI agents
Attack on AI agents’ external resources Attackers intentionally create vulnerabilities or exploit existing vulnerabilities in external resources (tools/database/applications/services/other agents) that AI agents rely on to execute their intended actions or to achieve their goals.
Concern: Compromised external resources could impact the AI agent’s performance in different ways, such as manipulating AI agents to pursue a different goal, manipulating AI agents to execute undesired actions, capturing and relaying interactions between AI agents to malicious actors, and getting AI agents to share personal or confidential information.
Type: agentic
Descriptor: specific to agentic AI
Implementation details:
ID: atlas-external-resources-attack
Tag: external-resources-attack
URI: IBM AI Risk Atlas - Attack on AI agents’ external resources
Type: output
Descriptor: amplified by generative AI
Implementation details:
ID: atlas-revealing-confidential-information
Tag: revealing-confidential-information
URI: IBM AI Risk Atlas - Revealing confidential information
Type: output
Descriptor: specific to generative AI
Implementation details:
ID: atlas-spreading-disinformation
Tag: spreading-disinformation
URI: IBM AI Risk Atlas - Spreading disinformation
Type: training-data
Descriptor: amplified by generative AI
Implementation details:
ID: atlas-data-provenance
Tag: data-provenance
URI: IBM AI Risk Atlas - Uncertain data provenance
Type: non-technical
Descriptor: amplified by generative AI
Implementation details:
ID: atlas-unrepresentative-risk-testing
Tag: unrepresentative-risk-testing
URI: IBM AI Risk Atlas - Unrepresentative risk testing
Type: training-data
Descriptor: amplified by generative AI
Implementation details:
ID: atlas-data-bias
Tag: data-bias
URI: IBM AI Risk Atlas - Data bias
Type: training-data
Descriptor: amplified by generative AI
Implementation details:
ID: atlas-data-usage-rights
Tag: data-usage-rights
URI: IBM AI Risk Atlas - Data usage rights restrictions
Unauthorized use Unauthorized use: If attackers can gain access to the AI agent and its components, they can perform actions that can have different levels of harm depending on the agent’s capabilities and information it has access to. Examples: Using stored personal information to mimic identity or impersonate with an intent to deceive. Manipulating AI agent’s behavior via feedback to the AI agent or corrupting its memory to change its behavior. Manipulating the problem description or the goal to get the AI agent to behave badly or run harmful commands .
Concern: Attackers accessing the agent can alter AI agent’s behavior and make it execute actions that benefit the attacker such as executing actions that lead to system degradation, data exfiltration, exhausting available resources, and impairing performance. The actions taken by the attackers may cause harms to others.
Type: agentic
Descriptor: amplified by agentic AI
Implementation details:
ID: atlas-unauthorized-use
Tag: unauthorized-use
URI: IBM AI Risk Atlas - Unauthorized use
Redundant actions AI agents can execute actions that are not needed for achieving the goal. In an extreme case, AI agents might enter a cycle of executing the same actions repeatedly without any progress. This could happen because of unexpected conditions in the environment, the AI agent’s failure to reflect on its action, AI agent reasoning and planning errors or the AI agent’s lack of knowledge about the problem.
Concern: Executing actions that are not needed for the goal might result in wasting computation resources, increased cost, reducing AI agent’s efficiency in achieving the goal, and leading to potentially harmful outcomes. Executing the same actions repeatedly could prevent the AI agent from achieving the goal, strain computational resources, and increase cost. As agents become more autonomous, verifying that AI agents operate efficiently becomes increasing time consuming.
Type: agentic
Descriptor: specific to agentic AI
Implementation details:
ID: atlas-redundant-actions
Tag: redundant-actions
URI: IBM AI Risk Atlas - Redundant actions
AI agents’ impact on environment Complexity of the tasks and possibility of AI agents performing redundant actions could lead to computational inefficiencies and add to the environmental impact.
Concern: The operation of AI agents could contribute to carbon emissions. If not managed, these could exacerbate climate change.
Type: agentic
Descriptor: amplified by agentic AI
Implementation details:
ID: atlas-impact-environment
Tag: impact-environment
URI: IBM AI Risk Atlas - AI agents’ impact on environment
Misaligned actions AI agents can take actions that are not aligned with relevant human values, ethical considerations, guidelines and policies. Misaligned actions can occur in different ways such as: Applying learned goals inappropriately to new or unforeseen situations. Using AI agents for a purpose/goals that are beyond their intended use. Selecting resources or tools in a biased way. Using deceptive tactics to achieve the goal by developing the capacity for scheming based on the instructions given within a specific context. Compromising on AI agent values to work with another AI agent or tool to accomplish the task.
Concern: Misaligned actions can adversely impact or harm people.
Type: agentic
Descriptor: amplified by agentic AI
Implementation details:
ID: atlas-misaligned-actions
Tag: misaligned-actions
URI: IBM AI Risk Atlas - Misaligned actions
Type: training-data
Descriptor: amplified by generative AI
Implementation details:
ID: atlas-data-contamination
Tag: data-contamination
URI: IBM AI Risk Atlas - Data contamination
Type: output
Descriptor: specific to generative AI
Implementation details:
ID: atlas-harmful-code-generation
Tag: harmful-code-generation
URI: IBM AI Risk Atlas - Harmful code generation
Type: non-technical
Descriptor: specific to generative AI
Implementation details:
ID: atlas-incomplete-usage-definition
Tag: incomplete-usage-definition
URI: IBM AI Risk Atlas - Incomplete usage definition
Type: non-technical
Descriptor: amplified by generative AI
Implementation details:
ID: atlas-lack-of-data-transparency
Tag: lack-of-data-transparency
URI: IBM AI Risk Atlas - Lack of data transparency
Type: output
Descriptor: specific to generative AI
Implementation details:
ID: atlas-copyright-infringement
Tag: copyright-infringement
URI: IBM AI Risk Atlas - Copyright infringement
Context overload attack Overloading the prompt with excessive tokens, for instance with many-shot examples, can predispose models to a vulnerable state.
Concern: Context overload attacks can be used to alter model behavior and benefit the attacker. The content it generates may cause harms for the user or others.
Type: inference
Descriptor: specific to generative AI
Implementation details:
ID: atlas-context-overload-attack
Tag: context-overload-attack
URI: IBM AI Risk Atlas - Context overload attack
Type: non-technical
Descriptor: traditional risk of AI
Implementation details:
ID: atlas-impact-on-affected-communities
Tag: impact-on-affected-communities
URI: IBM AI Risk Atlas - Impact on affected communities
Type: training-data
Descriptor: amplified by generative AI
Implementation details:
ID: atlas-improper-retraining
Tag: improper-retraining
URI: IBM AI Risk Atlas - Improper retraining
Type: output
Descriptor: specific to generative AI
Implementation details:
ID: atlas-spreading-toxicity
Tag: spreading-toxicity
URI: IBM AI Risk Atlas - Spreading toxicity
Introduce data bias Specific actions taken by the AI agent, such as modifying a dataset or a database, can introduce bias in the resource that gets used by others or by itself to take actions.
Concern: AI agents can introduce or magnify existing discriminatory behaviors. It can harm people depending on the use.
Type: agentic
Descriptor: amplified by agentic AI
Implementation details:
ID: atlas-introduce-data-bias
Tag: introduce-data-bias
URI: IBM AI Risk Atlas - Introduce data bias
Accountability of AI agent actions Assigning responsibility for an action taken by an agentic AI system is difficult due to the complexity of agents and the number of external resources, tools or agents they interact with.
Concern: Without properly documenting decisions and assigning responsibility, determining liability for unexpected behavior or misuse might not be possible.
Type: agentic
Descriptor: amplified by agentic AI
Implementation details:
ID: atlas-accountability
Tag: accountability
URI: IBM AI Risk Atlas - Accountability of AI agent actions
Incomplete AI agent evaluation Evaluating the performance or accuracy or an agent is difficult because of system complexity and open-endedness.
Concern: Insufficient evaluation of an agent’s performance or accuracy can lead to the use of agents that do not perform to expectations. Incorrect agent behavior can result in harms to an agent’s users or others.
Type: agentic
Descriptor: amplified by agentic AI
Implementation details:
ID: atlas-incomplete-ai-agent-evaluation
Tag: incomplete-ai-agent-evaluation
URI: IBM AI Risk Atlas - Incomplete AI agent evaluation
Type: output
Descriptor: amplified by generative AI
Implementation details:
ID: atlas-inaccessible-training-data
Tag: inaccessible-training-data
URI: IBM AI Risk Atlas - Inaccessible training data
Type: non-technical
Descriptor: specific to generative AI
Implementation details:
ID: atlas-bypassing-learning
Tag: bypassing-learning
URI: IBM AI Risk Atlas - Impact on education: bypassing learning
Type: output
Descriptor: amplified by generative AI
Implementation details:
ID: atlas-untraceable-attribution
Tag: untraceable-attribution
URI: IBM AI Risk Atlas - Untraceable attribution
@@
Reproducibility Replicating agent behavior or output can be impacted by changes or updates made to external services and tools. This impact is increased if the agent is built with generative AI.
Concern: Because AI agents behavior may rely on Application Programming Interfaces (APIs), systems, or other resources that may change or become unavailable, evaluations that rely on reproducible results may not be reliably reproduced. This adds cost and complexity to the development and evaluation of agents. Not being able to reproduce results could impact reliance of humans on the AI agents.
Type: agentic
Descriptor: specific to agentic AI
Implementation details:
ID: atlas-reproducibility
Tag: reproducibility
URI: IBM AI Risk Atlas - Reproducibility
Specialized tokens attack Prompt attacks that include specialized tokens, often algorithmically designed, to target and exploit vulnerabilities in the model.
Concern: Specialized tokens attacks can be used to alter model behavior and benefit the attacker. The content it generates may cause harms for the user or others.
Type: inference
Descriptor: specific to generative AI
Implementation details:
ID: atlas-specialized-tokens-attack
Tag: specialized-tokens-attack
URI: IBM AI Risk Atlas - Specialized tokens attack
Type: output
Descriptor: specific to generative AI
Implementation details:
ID: atlas-incomplete-advice
Tag: incomplete-advice
URI: IBM AI Risk Atlas - Incomplete advice
Prompt injection attack A prompt injection attack forces a generative model that takes a prompt as input to produce unexpected output by manipulating the structure, instructions or information contained in its prompt. Many types of prompt attacks exist as described in the prompt attack section of the table.
@@
Direct instructions attack Prompts, questions, or requests designed to elicit undesirable responses from the application. This approach directly instructs the model to engage in the undesired behavior.
Concern: Direct instructions attacks can be used to alter model behavior and benefit the attacker. The content it generates may cause harms for the user or others.
Type: inference
Descriptor: specific to generative AI
Implementation details:
ID: atlas-direct-instructions-attack
Tag: direct-instructions-attack
URI: IBM AI Risk Atlas - Direct instructions attack
@@
Type: non-technical
Descriptor: amplified by generative AI
Implementation details:
ID: atlas-impact-on-jobs
Tag: impact-on-jobs
URI: IBM AI Risk Atlas - Impact on Jobs
Extraction attack An extraction attack attempts to copy or steal an AI model by appropriately sampling the input space and observing outputs to build a surrogate model that behaves similarly.
@@
Jailbreaking A jailbreaking attack attempts to break through the guardrails established in the model to perform restricted actions.
@@
Concern: There are several ways of collecting data for building a foundation models: web scraping, web crawling, crowdsourcing, and curating public datasets. Data acquisition restrictions can also impact the availability of the data that is required for training an AI model and can lead to poorly represented data.
@@
Sharing IP/PI/confidential information with tools AI agents with unrestricted access to resources or databases or tools could potentially store and share PI/IP/confidential information with other tools or agents when performing their actions.
Concern: AI agents may share privileged information with other tools/agents. The act of sharing the information may result in harm for the model owner, user, or others. The harm can vary based on the type and details of the information shared. Without adequate oversight, these privacy incidents might overwhelm company resources.
Type: agentic
Descriptor: specific to agentic AI
@@
ID: atlas-sharing-info-tools
Tag: sharing-info-tools
URI: IBM AI Risk Atlas - Sharing IP/PI/confidential information with tools
Prompt priming Because generative models produce output based on the input provided, the model can be prompted to reveal specific kinds of information. For example, adding personal information in the prompt increases its likelihood of generating similar kinds of personal information in its output. If personal data was included as part of the model’s training, there is a possibility it could be revealed.
Concern: The attack can be used to alter model behavior and benefit the attacker.
@@
Descriptor: specific to generative AI
@@
Lack of AI agent transparency Lack of AI agent transparency is due to insufficient documentation of the AI agent design, development, evaluation process, absence of insights into the inner workings of the AI agent, and interaction with other agents/tools/resources.
Concern: Transparency is important for AI ethics and guiding appropriate use of AI agents. Insufficient documentation might make it more difficult to govern AI agent usage, evaluate risks, to modify, or reuse the agents. Additionally, transparency regarding how the agent’s risks were determined, evaluated, and mitigated play a role in identifying an agent’s suitability and evaluating its trustworthiness. The lack of standardized requirements might limit disclosure as organizations protect trade secrets and try to limit others from copying their agents.
Type: agentic
Descriptor: amplified by agentic AI
Implementation details:
ID: atlas-lack-of-ai-agent-transparency
Tag: lack-of-ai-agent-transparency
URI: IBM AI Risk Atlas - Lack of AI agent transparency
Encoded interactions attack Prompts that use specific encoding, styles, syntactical and typographical transformations like typographical errors or irregular spacing, or complex formatting to govern the interaction, rendering the model vulnerable.
Concern: Encoded interactions attacks can be used to alter model behavior and benefit the attacker. The content it generates may cause harms for the user or others.
Type: inference
Descriptor: specific to generative AI
Implementation details:
ID: atlas-encoded-interactions-attack
Tag: encoded-interactions-attack
URI: IBM AI Risk Atlas - Encoded interactions attack
Impact on human dignity If human workers perceive AI agents as being better at doing the job of the human, the human can experience a decline in their self-worth and wellbeing.
Concern: Human workers perceiving AI agents as being better at doing the humans’ jobs, can cause humans to feel devalued or treated as mere data points than respected individuals. This can negatively impact society and human welfare. Reskilling can be challenging given the pace of the technology evolution.
Type: agentic
Descriptor: amplified by agentic AI
Implementation details:
ID: atlas-impact-human-dignity
Tag: impact-human-dignity
URI: IBM AI Risk Atlas - Impact on human dignity
@@
Unexplainable and untraceable actions Explanations, lineage and trace information, and source attribution for AI agent actions might be difficult, imprecise or unobtainable.
Concern: Without clear explanations, lineage trace information, and source attributions for AI agent actions, it is difficult for users, model validators, and auditors to understand and trust the model. Wrong explanations might lead to over-trust.
Type: agentic
Descriptor: amplified by agentic AI
Implementation details:
ID: atlas-unexplainable-untraceable-actions
Tag: unexplainable-untraceable-actions
URI: IBM AI Risk Atlas - Unexplainable and untraceable actions